Compute for the strength of electric field if a certain charge is 2.5 µm away from a source charge that has a magnitude of 15 µC.
1
Expert's answer
2020-06-03T12:05:17-0400

The electric field strength


"E=k\frac{q}{r^2}"

"E=9\times 10^9\frac{15\times 10^{-6}}{(2.5\times 10^{-6})^2}=2.16\times 10^{16}\:\rm V/m"